For the given functions f and g, find (f times g)(x)
f(x)=7x+5, g(x)=3x-4
Can you please show me the steps to how to find the answer? I don't understand this problem at all

Answer by jim_thompson5910(35256)

(f * g)(x) = f(x) * g(x)


(f * g)(x) = (f(x)) * (g(x))


(f * g)(x) = (7x+5) * (3x-4)


(f * g)(x) = 7x * (3x-4) + 5 * (3x-4)


(f * g)(x) = 7x*3x + 7x*(-4) + 5*(3x) + 5*(-4)


(f * g)(x) = 21x^2 - 28x + 15x - 20


(f * g)(x) = 21x^2 - 13x - 20
------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------